OpenStudy (anonymous):

Find b and c so that y=-7x^2+bx+c has vertex (4,9). Can someone show me step by step? I got b=56 and c=-103

OpenStudy (amistre64):

hmm ... if we can turn this into the vertex form id be simpler i think

OpenStudy (amistre64):

y = -7x^2 + bx + c ; lets start by factoring out that -7 from the first 2 terms y = -7(x^2 - bx/7) + c ; how we can complete the square to change its form y = -7(x^2 - bx/7 + (2b/7)^2) + 7(2b/7)^2 + c now we can compact it ... maybe y = -7(x^2 - 2b/7)^2 + 7(2b/7)^2 + c this is the vertex form if i did it right

OpenStudy (amistre64):

i got my divide by 2 off; spose to be b/14, not 2b/7

OpenStudy (amistre64):

y = -7(x^2 - b/14)^2 + 7(b/14)^2 + c thats better, hopefully

OpenStudy (amistre64):

now, the vertex is: (b/14 , 7(b/14)^2 + c) which needs to equate to (4,9) b/14 = 4 7(b/14)^2 + c = 9 ----------------------------- b/14 = 4 b = 56 ......................... 7(b/14)^2 + c = 9 7(56/14)^2 + c = 9 c = 9 - 7(56/14)^2

OpenStudy (amistre64):

so yeah, b=56 and c=-103 :)

OpenStudy (anonymous):

thank you! i have been gone from my class on family emergency and have to learn this on my own. the book doesnt really help.

OpenStudy (amistre64):

i got no idea what your book would tell you to do, i just took what i knew and tried to keep me head on stratight till i got to the end :)

OpenStudy (anonymous):

haha yea!